# End-to-End Dexterous Arm-Hand VLA Policies via Shared Autonomy: VR Teleoperation Augmented by Autonomous Hand VLA Policy for Efficient Data Collection

### v1

@@ Abstract
Achieving human-like manipulation capabilities with dexterous hands for general-purpose robots remains a grand challenge. While recent Vision-Language-Action (VLA) models show promise in learning flexible skills from human-guided demonstrations, their scalability is constrained by the scarcity of high-quality training data. Existing real-robot data collection has certain inherent limitations: fully manual teleoperation imposes excessive cognitive load on human operators, limiting session duration, while automated planning often produces unnatural motions and yields a data distribution that is suboptimal for the learning targeted skillful manipulation.
To address this, we propose a Shared Autonomy framework that partitions control along the macro-micro motion domains. A human operator guides the pose of the robot end-effector via intuitive VR teleoperation, while an autonomous DexGrasp-VLA policy, using real-time tactile and local visual feedback, serves as an assistant for fine-grained and force-adaptive hand control. This division of labour significantly reduces human cognitive load and enables the efficient collection of high-quality data of coordinated arm-hand demonstrations with minimum mental fatigue.
Leveraging these demonstration data, we train an end-to-end VLA policy enhanced with our proposed novel Arm-Hand Feature Enhancement module. This architecture explicitly captures both the distinct latent features of macro (arm) and micro (hand) movements and their shared representations, resulting in more natural and robust arm-hand coordination. Furthermore, our Corrective Teleoperation system enables continuous policy improvement through human-in-the-loop failure recovery and data augmentation. Experiments show that our framework generates high-quality data with very low manpower requirements and the resulting fine-tuning can effectively learn policies that achieve around 90% success rate across a diverse set of over 50 objects, including unseen instances. The system’s effectiveness is validated through comprehensive evaluations and ablation studies, highlighting its potential to develop dexterous manipulation capabilities.
@@ 1 Introduction
To address these limitations, we propose a Shared Autonomy framework (Fig. 1) that partially offloads the high-level arm teleoperation from fine-grained low-level autonomous control of multiple fingers of the dexterous hand, enabling efficient collection of high-quality and coordinated demonstrations with low mental load. The core innovation lies in a functional division of shared control that leverages the complementary strengths of humans and AI (the autonomous VLA policy for the hand): the human focuses on high-level scene understanding, semantics, spatial reasoning, and affordance to teleoperate the robot end-effector easily via the Virtual Reality (VR) interface. Meanwhile, a learned DexGrasp-VLA controller acts as a VLA Copilot that autonomously executes fine-grained and compliant grasping motions with the dexterous hand. The DexGrasp-VLA copilot for the hand is a tactile-based VLA policy for the dexterous hand. It integrates language commands, in-hand camera images, proprioceptive states, and two complementary tactile features: resultant force vectors (magnitude and direction) and spatial tactile embeddings (contact distribution patterns). This multi-modal perception provides sufficient feedback for robust and adaptive grasping performance across diverse conditions. This shared-autonomy paradigm eliminates the need for the operator to simultaneously control all degrees of freedom of a complex arm-hand system. By delegating high-DoF hand control to the VLA Copilot, we significantly reduce the cognitive load from the human operator, produce naturally coordinated arm-hand trajectories with stable grasps, and enable faster data collection with a high success rate.
Figure 1: Fully manual teleoperation versus the proposed Shared Autonomy – semi-autonomous teleoperation, where the human operator focuses on high-level control of the arm motion, while the hand is controlled by the perceptual AI-Copilot for autonomous grasping.
@@
Our primary contributions are as follows:
  • Multimodal VLA Copilot for Dexterous Grasping: a VLA Copilot, DexGrasp-VLA, that fuses visual, language, tactile, and proprioceptive feedback to autonomously perform force-adaptive grasping for a five-finger hand.
  • Shared Autonomy for Data Collection: human VR teleoperation combined with a VLA Copilot for autonomous grasping, enabling efficient collection of high-quality demonstrations with low cognitive load.
  • End-to-End VLA with Arm-Hand Feature Enhancement: an end-to-end VLA policy for the holistic arm-hand control and coordination using a novel architecture with distinct arm and hand features plus shared representations, capturing both macro- and micro-motions for robust and natural control.
  • Corrective Human-in-the-Loop Teleoperation: a continuous learning strategy that incorporates both successful trajectories plus recovery data during deployment for iterative policy refinement.
@@
Our experimental results demonstrate the effectiveness of the proposed framework. It enables the efficient data collection of high-quality and coordinated arm-hand demonstrations, which in turn underpins the learning of an end-to-end VLA policy that achieves a 90% success rate across a diverse set of 50 more objects, including unseen objects and settings. The necessity of each core component was systematically validated through ablation studies, which confirmed that the DexGrasp-VLA model, the Arm-Hand Feature Enhancement module, and the Corrective Teleoperation system are all essential and significantly improve the policy’s performance, success rate, and robustness.
@@ 2.1 Vision-Language-Action Models for Robot Control
An increasing amount of research work has been investigating how to extend the VLA frameworks to dexterous hand manipulation. The work in [56] proposes its own DexGraspVLA model as a hierarchical architecture that combines a pre-trained vision-language planner with a diffusion-based low-level action controller. It transforms diverse language and visual inputs into domain-invariant representations, enabling robust imitation learning. Being-H0 [34] takes a data-centric approach, pretraining VLA models on large-scale human hand videos with a novel physical instruction tuning paradigm. The model utilizes part-level motion tokenization for millimeter-accurate trajectory learning and 3D physical space alignment, thereby enhancing generalization. In parallel, other research has been conducted to incorporate tactile sensing into VLA models. For example, OmniVTLA [12] introduces a Vision-Tactile-Language-Action architecture with a dual-path tactile encoder, consisting of a vision transformer and a semantically-aligned tactile ViT.
@@
Based on the first principle of the different characteristics in motion and physical interactions between the hand and arm, this monolithic design in VLA leads to two key limitations. First, training the arm-hand policies in a monolithic manner fails to capture the distinct kinematic and dynamic properties of macro-level arm motion versus micro-level hand manipulation, potentially limiting coordination and robustness. Second, these systems rely heavily on large-scale, offline-collected demonstrations, either through scripted control policies or direct full-DoF teleoperation [36] with exoskeletons, which are costly and difficult to scale, creating a data bottleneck especially for dexterous hands. Moreover, these systems often employ an undifferentiated fusion of tactile features, lacking explicit representation of contact information for more precise action generation of the VLAs to fine control hand motions.
@@ 2.3 Tactile Sensing for Robot Manipulation
Tactile sensing provides essential feedback for capturing contact dynamics during physical interactions, especially in tasks involving compliant interaction, precision, or uncertainty. Prior works have explored vision-tactile fusion in reinforcement learning, demonstrating effectiveness in assembly [18, 27] and dexterous in-hand manipulation [33, 19]. These methods typically use tactile signals as feedback for closed-loop policies in contact-rich scenarios. More recently, an increasing number of research apply imitation learning to learn combined vision-tactile representations for fine-grained control [51, 30, 20, 28]. These models enable reactive manipulation using end-to-end learning from multimodal demonstrations.
Building on these previous works, recent methods have begun to incorporate tactile perception into Vision-Language-Action (VLA) frameworks [3, 21, 12, 53], improving generalization and robustness across object categories and tasks. For example, VLA-Touch [3] enhances a frozen VLA backbone by introducing dual-level tactile feedback: a pretrained tactile-language model assists in high-level semantic planning, and a diffusion controller refines low-level execution. Tactile-VLA [21] further integrates tactile signals into both planning and control, combining a hybrid position–force controller with commonsense reasoning based on vision-language priors. This design enables zero-shot generalization in contact-rich tasks using minimal tactile supervision. VTLA [53] focuses on fine contact manipulation, such as insertion, and proposes a Vision-Tactile-Language model trained via Direct Preference Optimization. It achieves high success rates in both simulation and real-world peg-in-hole tasks. OmniVTLA [12] takes a more data-centric approach by constructing a large tri-modal dataset and learning semantically aligned tactile embeddings across diverse sensor modalities, leading to strongly improved performance for both the gripper and dexterous hand.
Despite of the nuanced differences in technical details, these methods all attempt to fuse tactile inputs into the VLA backbone, either through full end-to-end retraining, multimodal pretraining, or task-specific architectural improvements. Such strategies often rely on large-scale tactile datasets, or require architectural changes that add significant complexity to the model. In contrast, our method opts for a more modular design approach: we inject tactile feedback locally into the action expert and enrich the contact-relevant information to improve contact-rich control, while keeping the vision-language backbone totally intact which introduces no complexity compared to some of the aforementioned methods. This allows for efficient learning and deployment without full model retraining, and still preserves the usefulness of tactile grounding during execution.
@@ 3 Methods and Training Recipe
  • Second, within the shared autonomy framework, a human operator teleoperates the arm via a VR interface, while the DexGrasp-VLA acts as a VLA Copilot for controlling the hand. This division of labour enables efficient collection of synchronized arm-hand demonstration data with significantly reduced cognitive load for the operator.
  • Third, using the collected demonstration data, we train a holistic end-to-end policy augmented with our novel arm-hand feature enhancement module. This architecture extracts and fuses dedicated features specifically for the arm and hand, which significantly improves the coordination between macro- and micro- movements.
@@ 3.2 DexGrasp-VLA: Autonomous Dexterous Grasping Policy
Figure 3: Data collection and training framework for force-adaptive grasping policy learned via LSTM. The datasets are collected by parameterized control (68 demos) and teleoperated human strategies (150 demos), respectively: the former captures the implicit tactile/force-sensitive grasping, and the latter incorporates more flexible multi-finger form closure from human teleoperation, ensuring data diversity.
@@ 3.2.1 Force-Adaptive Grasping Policy Learned by LSTM
To bootstrap the learning of reactive grasping under rich contact dynamics without relying on complex visual perception, we first train a “blind” policy [49] that depends solely on proprioceptive and tactile feedback. The policy is trained using a hybrid dataset combining trajectories from a parametric force-based controller (68 demos) and human teleoperation demonstrations (150 demos), ensuring both physical compliance and human-like grasping strategies. We use a diverse set of over 20 objects with varying shapes, sizes, weights, and surface materials to enhance the implicit behavior of variable adaptive grip forces, given the history information of proprioceptive and tactile sensing encapsulated by the LSTM.
@@
Data are collected through two complementary methods: 1) Force-Adaptive Position Control: A rule-based controller generates compliant grasping motions at 50 Hz. At each step $i$, the command $q_{c}(i)$ is computed as:
@@
where $q_{m}(i)$ is the measured joint position, $q(0)$ is the initial position, $k$ is the gain to adjust grasping force , and $f_{z}(i)$ is the resultant normal force at the fingertips. This control law enables rapid hand closure when no contact is detected and gradually increases the grip force upon contact, emulating human-like force modulation for stable grasping.
2) Human Teleoperation via Retargeting: We use a Leap Motion sensor to capture human hand motions, which are retargeted to the 12-DoF Xhand. Each demonstration includes the full grasping sequence from hand opening to stable grasp, sampled at 50 Hz.
@@
To unify the diverse strategies from both automated and teleoperation-based data collection pipelines, we train a lightweight LSTM-based policy (Fig. 3) that fuses these approaches into a compact state-based controller. This design achieves three primary objectives: (i) computational efficiency for real-time deployment on physical hardware, (ii) generalization across diverse object shapes, materials, and contact conditions by capturing temporal dependencies in tactile-proprioceptive sequences, and (iii) scalable autonomous data generation without requiring continuous human supervision. Once trained, this policy serves as both a unified controller bridging heterogeneous data sources and a practical expert for bootstrapping large-scale training of vision-based VLA policies.
We implement this approach through behavior cloning, where the input at each time step $t$ is represented as $x_{t}=[s_{t}^{\text{hand}},f_{t}^{\text{hand}}]\in\mathbb{R}^{39}$, with $s_{t}=[q_{t}^{\text{hand}},\tau_{t}^{\text{hand}}]\in\mathbb{R}^{24}$ denoting the proprioceptive state and $f_{t}^{\text{hand}}\in\mathbb{R}^{15}$ representing contact force signals measured from the fingertip tactile arrays. A sliding window of length $T$ constructs the sequence input $X=[x_{t-T+1},\dots,x_{t}]\in\mathbb{R}^{T\times 39}$, ensuring temporal continuity.
The policy network uses two parallel three-layer MLPs with ReLU activations and batch normalization to independently encode proprioceptive and tactile inputs. Their outputs are concatenated to form a fused feature representation $H\in\mathbb{R}^{T\times 256}$, which is passed into a single-layer LSTM with the hidden layer size of 256 to capture temporal dependencies critical for contact-rich manipulation. The output at the final time step is mapped through a fully connected layer to predict the hand action $a_{t}\in\mathbb{R}^{12}$.
The model is optimized with a mean squared error (MSE) loss between predicted and demonstrated actions, regularized with an $\mathcal{L}_{2}$ penalty to prevent overfitting:
@@
The trained LSTM policy functions as a vision-free reactive expert, producing robust grasping trajectories from tactile feedback. Its main role is to enable large-scale autonomous data collection, thereby enriching training diversity and improving the generalization of the subsequent VLA policy. It should be noted that different object properties such as stiffness defamation and so will be indirectly captured by history information of proprioceptive and tactile sensing, in which the diverse multi-finger form-closure is provided by the human demonstration data, and steady and appropriate grip force control – without damaging or crushing the object – is provided by force-control based method where the settle-down force is implicitly regulated as in Equation 1 in an admittance manner. The future generation of force controller dexterity hands can achieve this force-adaptive behavior more easily by directly setting reference contact forces suitable for various objects, which could achieve much higher performance.

### v2

@@ Abstract
Dexterous manipulation with multi-finger robotic hands remains a significant challenge for general-purpose robots. Recent Vision-Language-Action (VLA) models show promise in learning flexible skills from human-guided demonstrations, yet their scalability is limited by the scarcity of high-quality training data. Existing real-robot data collection faces inherent constraints: fully manual teleoperation imposes excessive cognitive load on human operators, limiting session duration, while automated planning often produces stiff, unnatural motions and yields suboptimal data distribution for learning skillful manipulation.
To address this, we propose a Shared Autonomy framework that partitions control along the macro-micro motion domains. A human operator guides the robot end-effector via intuitive VR teleoperation, while an autonomous DexGrasp-VLA policy, using real-time tactile and local visual feedback, assists fine-grained and force-adaptive hand control as a Copilot. This division of labor significantly reduces human cognitive load and enables efficient collection of high-quality data of coordinated arm-hand demonstrations with minimum mental fatigue.
Using these demonstration data, we train an end-to-end VLA policy enhanced with a novel Arm-Hand Feature Enhancement module. This architecture explicitly captures both the distinct latent features of macro (arm) and micro (hand) movements and their shared representations, resulting in smoother and more robust arm-hand coordination. Furthermore, our Corrective Teleoperation system supports continuous policy improvement through human-in-the-loop failure recovery and data augmentation. Experiments show that the framework produces high-quality data with very low operator effort, and the resulting fine-tuning can effectively train policies that achieve around 90% success rate across more than 50 diverse objects, including unseen instances. The system’s effectiveness is validated through comprehensive evaluations and ablation studies, highlighting its potential for advancing dexterous robotic manipulation.
@@ 1 Introduction
To address these limitations, we propose a Shared Autonomy framework (Fig. 1) that partially offloads the high-level arm teleoperation from fine-grained low-level autonomous control of multiple fingers of the dexterous hand, enabling efficient collection of high-quality and coordinated demonstrations with low mental load. The core innovation lies in a functional division of shared control that leverages the complementary strengths of humans and AI (the autonomous VLA policy for the hand): the human focuses on high-level scene understanding, semantics, spatial reasoning, and affordance to teleoperate the robot end-effector easily via the Virtual Reality (VR) interface. Meanwhile, a learned DexGrasp-VLA controller acts as a VLA Copilot that autonomously executes fine-grained and compliant grasping motions with the dexterous hand. The DexGrasp-VLA copilot for the hand is a tactile-based VLA policy for the dexterous hand. It integrates language commands, in-hand camera images, proprioceptive states, and two complementary tactile features: resultant force vectors (magnitude and direction) and spatial tactile embeddings (contact distribution patterns). This multi-modal perception of the DexGrasp-VLA provides sufficient feedback for robust and adaptive grasping performance across diverse conditions.
This shared-autonomy paradigm eliminates the need for the operator to simultaneously control all degrees of freedom of a complex arm-hand system. By delegating high-DoF hand control to the VLA Copilot, we significantly reduce the cognitive load from the human operator, produce naturally coordinated arm-hand trajectories with stable grasps, and enable faster data collection with a high success rate. It shall be noted that this work uses VLA policies in two different ways: one is the hand-only tactile-aware VLA, and the other is for the holistic arm-hand control – the DexGrasp-VLA is a sub-policy for controlling the dexterous hand within the full VLA framework for data collection and model training.
Figure 1: Comparison of fully manual teleoperation and the proposed Shared Autonomy – semi-autonomous teleoperation, where the human operator focuses on high-level control of the arm motion, while the hand is controlled by the perceptual AI-Copilot for autonomous grasping.
@@
The core technical contributions are as follows:
  • Novel Multimodal VLA Copilot for Dexterous Grasping: We present DexGrasp-VLA, the first Vision-Language-Action (VLA) Copilot that robustly fuses visual, tactile, language, and proprioceptive feedback for autonomous, force-adaptive grasping with a five-finger hand. Its reactive and compliant fine-grained grasping capability represents a major advance over prior data collection approaches which lacked consistent performance in contact handling and force adaptability.
  • Shared Autonomy for Efficient Data Collection: We propose a paradigm-shifting framework that strategically divides control between the human operator and the VLA Copilot to overcome the data bottleneck in dexterous manipulation. By combining intuitive human VR teleoperation for high-level arm guidance assisted with the tactile-enhanced DexGrasp-VLA Copilot for autonomous fine-grained grasping, we substantially reduce operator cognitive load and enable efficient collection of high-quality demonstration data.
  • End-to-End VLA with Arm-Hand Feature Enhancement: We design a novel VLA architecture that explicitly disentangles the control of macro (arm) and micro (hand) movements through dedicated feature pathways, while preserving a shared global task representation. This architectural innovation directly solves the problem of monolithic controllers that fail to capture the distinct kinematics and dynamics of the arm and hand respectively, thus leading to significantly more robust and generalizable arm-hand coordination, with notably improved resilience to visual occlusions.
  • Corrective Human-in-the-Loop Teleoperation for Continuous Policy Improvement: We introduce a corrective human-in-the-loop intervention that operates without interrupting deployed policy execution, enabling purposeful data collection from long-tail failure cases during real-world deployment. Such teleoperation allows humans to assist imperfect VLA executions, producing complementary datasets for scenarios that would otherwise fail. Integrating these curated recovery data with successful trajectories, we establish a self-improving data flywheel: retraining the base VLA model on the iteratively augmented dataset results in an updated policy with enhanced robustness in real-world applications. This approach extends the training by unifying both pre-scripted demonstrations with in-situ deployment data, thereby enabling targeted coverage of corner cases that are otherwise impractical to anticipate beforehand.
@@
Our experimental results demonstrate the effectiveness of the proposed framework. It enables the efficient data collection of high-quality and coordinated arm-hand demonstrations, which in turn underpins the learning of an end-to-end VLA policy that achieves a 90% success rate across a diverse set of over 50 objects, including unseen objects and settings. The necessity of each core component was systematically validated through comprehensive ablation studies, which confirmed that the DexGrasp-VLA model, the Arm-Hand Feature Enhancement module, and the Corrective Teleoperation system are all essential and significantly improve the policy’s performance, success rate, and robustness. These strong empirical results provide an essential testbed for exploring broader algorithmic strategies before scaling up to larger models.
This work is strategically positioned within the broader challenge of developing robust foundational models for embodied intelligence with a focus on dexterous manipulation. While scaling laws have achieved remarkable progress in large vision and language models, the direct application to robotics still faces significant data quality and architectural challenges, due to the fact that motion trajectories, visual streams, and high-frequency tactile-force signals are fundamentally heterogeneous in nature. Simply brute-force directing heterogeneous data into a monolithic architecture is unlikely to yield the desired emergent dexterity, or effectively leverage scaling laws in a scientific manner.
Our work argues that a necessary premise to effective large-scale pre-training is the foundational understanding of how to algorithmically integrate these distinct sensorimotor pathways. By leveraging SFT of pre-trained VLA models, we agilely iterate research investigation: we can rapidly isolate and validate core components, such as our Arm-Hand Feature Enhancement module and effective tactile integration strategies, within a controlled setting. This approach efficiently exploits the pre-trained models’ semantic and generalization capabilities, and meanwhile allowing us to focus on the fundamental physical principles of useful multimodal fusion. Thus, the effective algorithms developed and validated on a smaller scale provide the essential blueprint for the future collection of massive, high-quality datasets (e.g., via our Shared Autonomy framework), and strengthen higher likelihood of success in the subsequent pre-training of more sophisticated large-scale robot foundational models with true human-like physical intelligence.
@@ 2.1 Vision-Language-Action Models for Robot Control
An increasing amount of research work has been investigating how to extend the VLA frameworks to dexterous hand manipulation. The work in [56] proposes a DexGraspVLA model as a hierarchical architecture that combines a pre-trained vision-language planner with a diffusion-based low-level action controller. It transforms diverse language and visual inputs into domain-invariant representations, enabling robust imitation learning. Being-H0 [34] takes a data-centric approach, pretraining VLA models on large-scale human hand videos with a novel physical instruction tuning paradigm. The model utilizes part-level motion tokenization for millimeter-accurate trajectory learning and 3D physical space alignment, thereby enhancing generalization. In parallel, other research has been conducted to incorporate tactile sensing into VLA models. For example, OmniVTLA [12] introduces a Vision-Tactile-Language-Action architecture with a dual-path tactile encoder, consisting of a vision transformer and a semantically-aligned tactile ViT.
@@
Based on the first principle of the different characteristics in motion and physical interactions between the hand and arm, a monolithic design in VLA leads to two key limitations. First, training the arm-hand policies in a monolithic manner fails to capture the distinct kinematic and dynamic properties of macro-level arm motion versus micro-level hand manipulation, potentially limiting coordination and robustness. Second, these systems rely heavily on large-scale, offline-collected demonstrations, either through scripted control policies or direct full-DoF teleoperation [36] with exoskeletons, which are costly and difficult to scale, creating a data bottleneck especially for dexterous hands. Moreover, these systems often employ an undifferentiated fusion of tactile features, lacking explicit representation of contact information for more precise action generation of the VLAs to fine control hand motions.
@@ 2.3 Tactile Sensing for Robot Manipulation
Tactile sensing provides essential feedback for capturing contact dynamics during physical interactions, especially in tasks involving compliant interaction, precision, or uncertainty. Prior works have explored vision-tactile fusion in reinforcement learning, demonstrating effectiveness in assembly [18, 27] and dexterous in-hand manipulation [33, 19]. These methods typically use tactile signals as feedback for closed-loop policies in contact-rich scenarios. More recently, studies applied imitation learning to learn combined vision-tactile representations for fine-grained control [51, 30, 20, 28], enabling reactive manipulation with end-to-end learning from multimodal demonstrations.
Building on these, recent methods begin to incorporate tactile perception into Vision-Language-Action (VLA) frameworks [3, 21, 12, 53], improving generalization and robustness across object categories and tasks. For example, VLA-Touch [3] enhances a frozen VLA backbone with dual-level tactile feedback: a pretrained tactile-language model assists in high-level semantic planning, and a diffusion controller refines low-level execution. Tactile-VLA [21] further integrates tactile signals into planning and control, combining hybrid position–force control with commonsense reasoning based on vision-language priors. This design enables zero-shot generalization using minimal tactile supervision. VTLA [53] focuses on fine contact tasks such as insertion, proposing a Vision-Tactile-Language model trained via Direct Preference Optimization, achieving high success in both simulation and real-world peg-in-hole tasks. OmniVTLA [12] adopts a data-centric approach, constructing a large tri-modal dataset and learning semantically aligned tactile embeddings across diverse sensors, improving performance for both grippers and dexterous hands.
Despite nuanced differences in technical details, these methods all attempt to fuse tactile inputs into the VLA backbone, through full end-to-end retraining, multimodal pretraining, or task-specific architectural improvements. Such strategies often rely on large-scale tactile datasets, or require architectural changes that add significant complexity to the model. In contrast, our method uses a modular design approach: we inject tactile feedback locally into the action expert to enrich the contact-relevant information for improving contact-rich control, while keeping the vision-language backbone totally intact, adding no extra complexity. This allows for efficient learning and deployment without full retraining while preserving useful tactile grounding during execution.
@@ 3 Methods and Training Recipe
  • Second, within the shared autonomy framework, a human operator teleoperates the arm via a VR interface, while the DexGrasp-VLA acts as a VLA Copilot for controlling the hand. This division of labor enables efficient collection of synchronized arm-hand data with significantly reduced cognitive load for the operator.
  • Third, using such synchronized data, we train a holistic end-to-end policy augmented with our novel arm-hand feature enhancement module. This architecture extracts and fuses dedicated features specifically for the arm and hand, which significantly improves the coordination between macro- and micro- movements.
@@ 3.2.1 Force-Adaptive Grasping Policy Learned by LSTM
Figure 3: Data collection and training framework for the force-adaptive grasping policy learned via LSTM. The datasets are collected by parameterized control (68 demos) and teleoperated human strategies (150 demos), respectively: the former captures the implicit tactile/force-sensitive grasping, and the latter incorporates more flexible multi-finger form closure from human teleoperation, ensuring data diversity.
To bootstrap reactive grasping under rich contact dynamics without complex visual perception, we first train a “blind” policy [49] using solely the proprioceptive and tactile feedback. The policy uses a hybrid dataset from a parametric force-based controller (68 demos) and human teleoperation demonstrations (150 demos), ensuring force-sensitive compliance control and human-like grasping strategies. We use a diverse set of over 20 objects with varying shapes, sizes, weights, and materials to enhance the implicit behavior of variable adaptive grip forces, given the proprioceptive–tactile history encapsulated by the LSTM.
@@
Data are collected through two complementary methods: 1) Force-Adaptive Position Control: A rule-based controller generates compliant grasping at 50 Hz. At each step $i$, the command $q_{c}(i)$ is computed as:
@@
where $q_{m}(i)$ is the measured joint position, $q(0)$ is the initial position, $k$ is the gain to adjust the grasping force , and $f_{z}(i)$ is the fingertip resultant normal force. This control law enables rapid hand closure when no contact is detected, and gradually increases the grip force upon contact.
2) Human Teleoperation via Retargeting: A Leap Motion sensor is used to capture human hand motions, retargeted to the 12-DoF Xhand. Each demonstration covers the full grasping sequence from hand opening to stable grasp, sampled at 50 Hz.
@@
As shown in Fig. 3, unifying the diverse behaviors from automated and teleoperation-based data, we train a lightweight LSTM-based policy that fuses these grasping strategies into a compact state-based controller. This design achieves three objectives: (i) computational efficiency for real-time onboard deployment, (ii) generalization across diverse objects and contact conditions by capturing temporal dependencies in tactile-proprioceptive history, and (iii) scalable autonomous data generation without continuous human supervision. Once trained, this policy serves as a unified controller bridging heterogeneous data sources and a practical control expert to bootstrap training of vision-based VLA policies.
We implement through behavior cloning, where the input at each time step $t$ is represented as $x_{t}=[s_{t}^{\text{hand}},f_{t}^{\text{hand}}]\in\mathbb{R}^{39}$, with $s_{t}=[q_{t}^{\text{hand}},\tau_{t}^{\text{hand}}]\in\mathbb{R}^{24}$ denoting the proprioceptive state, and $f_{t}^{\text{hand}}\in\mathbb{R}^{15}$ representing signals measured from fingertip tactile arrays. A sliding window of length $T$ forms the sequence input $X=[x_{t-T+1},\dots,x_{t}]\in\mathbb{R}^{T\times 39}$, ensuring temporal continuity.
The policy network uses two parallel three-layer MLPs with ReLU and batch normalization to encode proprioceptive and tactile inputs independently. Outputs are concatenated into a fused feature representation $H\in\mathbb{R}^{T\times 256}$, passed into a single-layer LSTM (hidden layer size of 256) to capture essential temporal dependencies for contact-rich interactions. The output is mapped through a fully connected layer to predict the hand action $a_{t}\in\mathbb{R}^{12}$.
The model is optimized with a mean squared error (MSE) loss between predicted and demonstrated actions, plus a $\mathcal{L}_{2}$ regularization:
@@
The trained LSTM policy functions as a vision-free reactive expert, producing robust grasping trajectories from tactile feedback. Its main role is to make large-scale autonomous data collection easier, thereby enriching data diversity and improving VLA policy generalization. It should be noted that different object properties such as stiffness are indirectly captured by proprioceptive–tactile history; diverse multi-finger closures stem from human demos, while stable grip force and compliance control (without damaging or crushing objects) are provided by the force-control method, where the settle-down force is implicitly learned by behavior cloning in an admittance manner (Equation 1). Future dexterous hands with advanced force control could achieve this behavior directly by controlling reference contact forces for different objects, offering higher performance.